OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. R(0, ), S(-4, 2) m = 5

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}\] \[5=\frac{2-a}{-4-0}\] -20=2-a a=22